OpenStudy (anonymous):

PLEASE HELP!!! ILL GIVE MEDAL!!!!! Use the distributive property to rewrite the expression and solve. 9 x 72 = 9 x ???? + 9 x 2 = ???????

OpenStudy (anonymous):

The distributive property looks like this: 2(a + b) = 2a + 2b. Make sense?

OpenStudy (anonymous):

@AnyTipzWouldHelp can u plz come and help me plz ill right u a testimonal

OpenStudy (anonymous):

Dark, don't post on others' questions and it not be helping them, and no, you just want answers, not help, and that violates honor code.

OpenStudy (anonymous):

ya n i want to know how to do this so i dont have to always ask for help with it

OpenStudy (anonymous):

sorry and i do actuall want help i would like for u to explain how u got it i write that stuff down in my math book @AnyTipzWouldHelp

OpenStudy (anonymous):

I'll look at it after this dark, but one sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok kayla, look for common terms. You have a 9 in both equations, so you can use 9(72 + 2) instead of having two seperate equations.

OpenStudy (anonymous):

So the re-written equation is 9(72 + 2) = ????? 648 + 18 = ????? 666 = ????? That is your answer.

OpenStudy (anonymous):

@AnyTipzWouldHelp thank you :)

OpenStudy (anonymous):

No problem :)